# Score Distillation Sampling with Learned Manifold Corrective

### v1

@@ 5.2 Image-to-image Translation Network Training
Optimization-based editing is not always an optimal choice, e.g. when building interactive editing tools where execution speed matters. Further, optimization can be unstable or perform better or worse, depending on the input images or target prompts. To this end, we explore to what extend LMC-SDS can be used to train a reliable and fast image-to-image translation network. Given a known distribution of images, we train a network to translate images from that distribution to an unseen target distribution, only described by a target caption. Inspired by the authors of DDS [8], we train a cats-to-others network, with “others” being bears, dogs, lions, and squirrels. The network transforms an image of a cat into an image of one of the other species. In DDS, the authors fine-tune a pre-trained diffusion model into a single-step feed forward model using images created by that same diffusion model. In contrast, we train a standard U-Net from scratch and use a dataset of real images, a much more practical and versatile approach. Concretely, we train using cat images from AFHQ [3] and evaluate on 5% held out test images. We use the target prompt “a photo of a [class]” with $\omega=15$ for all methods and the source prompt “a photo of a cat” for DDS. Moreover, we stabilize the training using an L2 loss on the original image and use Cosine scheduling to let the network first focus on reconstructing the original image and then on edits, similarly to the experiment in DDS. The class is passed via a frequency encoded class id concatenated to each pixel. We report numerical results in Tab. 1 and show qualitative examples in Fig. 7. SDS performs similarly to previous experiments for this task. While the target class is evident in the resulting images, the results feature strong deficiencies. Again DDS struggles with significant edits and stays close to the source. In contrast, the network trained with our LMC-SDS loss produces realistic and convincing edits. This is also evident from the significantly higher CLIP-R-Precision among all network variants.
Figure 8: Results of DreamFusion using the original loss formulation (right) and our proposed LMC-SDS (left) for $\omega=20$. Our results are much sharper and contain more detail in all examples.
@@ 5.3 Text-to-3D Asset Generation
In our final experiment, we demonstrate how LMC-SDS can improve text-to-3D models in the style of DreamFusion. In Sec. 3, we discussed how large guidance weights $\omega$ may affect the quality of results. To provide further evidence for this observation, we produce 3D assets with varying $\omega$ using the original DreamFusion formulation and a variant using our LMC-SDS. The results in Fig. 9 also show the trend of high $\omega$ having the tendency to produce oversaturated colors, albeit not as as strongly as in the original paper, possibly due the use of a different image diffusion model. But we also observed an additional issue: the infamous Janus problem, the tendency to repeat detail in each view, becomes more severe with bigger $\omega$, as evident from the eagle example. On the other hand, when using low $\omega$ in the original SDS loss, we observe that DreamFusion struggles to create detailed structure or even structure at all. In contrast, when using LMC-SDS, DreamFusion successfully produces detailed results, even when using up to only one tenth of the originally proposed $\omega$. In Fig. 8, we show additional side-by-side comparisons of the original SDS-based DreamFusion and with LMC-SDS. Using our loss, the results are consistently sharper and feature far more detail. Finally in Fig. 10, we show that fixed $\boldsymbol{\epsilon}$ sampling may also be used in 3D asset generation, to obtain diverse results.
@@ 6 Discussion & Conclusion
We proposed a novel loss formulation called LMC-SDS. LMC-SDS allows to effectively utilize text-to-image diffusion models as rich image priors in several optimization and network training tasks. We derived LMC-SDS by analyzing DreamFusion’s SDS loss and by identifying an inherent problem in its formulation. Concretely, we expressed the SDS gradient as the sum of two terms. We identified that the second term $\mathcal{L}_{\text{proj}}$ provides deficient gradients and consequently was weighted down in DreamFusion. We argued that this term, however, should provide gradients towards the learned manifold of natural images associated to the diffusion model. We also provided evidence that the lack thereof makes the optimization unstable, thus negatively impacting the fidelity of results. To this end, we proposed to model the denoising models’ time step dependent image corruption. This allowed us to factor these out and obtain the cleaner gradient $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{lmc}}$. In extensive experiments, we have demonstrated the effectiveness and versatility of our novel LMC-SDS loss. Furthermore, we discussed how LMC-SDS differs from other SDS-style losses and demonstrated how its performance is superior both qualitatively and quantitatively.
@@
LMC-SDS finds its limits where either the signal from the diffusion model is too weak and ambiguous or when the optimization goes off the natural image manifold by too much. Concretely, LMC-SDS can only utilize whatever signal is coming from the diffusion model. This means, that when the diffusion model does not “understand” the prompt, LMC-SDS (and all other SDS-based approaches) will not be able to provide meaningful gradients. Secondly, whenever the current optimization state is too corrupted, LMC-SDS will not be able to correct for that and guide the optimization back to the manifold. Nevertheless, LMC-SDS is an important first step towards more stable and meaningful usage of the inherent image prior associated to image diffusion models.
We proposed LMC-SDS, a novel SDS-based loss formulation with cleaner gradient towards the image manifold. In the future, we would like to continue to investigate how the manifold corrective can be further improved. We also want to leverage our findings in concrete applications like text-to-3D or in image editing and stylization tasks.
@@
Appendix A Implementation Details
@@
We trained a standard U-Net with eleven layers with two Conv/ReLU/MaxPool blocks per layer, and skip connections from the encoder to the decoder. The first layer contains 32 filters and the number of filters is doubled until 256 in each layer in the encoder and halved in last layers of the decoder. The class id is a float value encoded using 10-dimensional frequency encoding and concatenated to each pixel. The final Conv layer with $\tanh$ activation outputs images of $256\times 256$px resolution. We trained the network with a learning rate of $5\times 10^{-5}$ with a decay of $0.9$ every $50$K steps. In the first $20$K steps, we use Cosine scheduling to increase the weight of $\mathcal{L}_{\text{LMC-SDS}}$ from $0.1$ to $1.0$ and decrease the weight of L2 regularization w.r.t. the original image from $1.0$ to $0.1$. We trained for $200$K iterations with batches of $32$.
@@
Appendix B Additional Results
Figure 11: Progression of an image optimized using the prompts “photo of a young girl” and “photo of an old woman”.
@@ Appendix B Additional Results
In Figs. 15 and 16 we present additional editing results. In contrast to the main paper, we show results for each prompt on a number of images. Concretely, we synthesized photographs of people and edited each of them using prompts describing different art styles or materials. Each result respects the corresponding prompt well and yet has an individual touch to it. The image statistics are well preserved and artifacts are rare. The overall quality once more demonstrates the effectiveness of our novel LMC-SDS loss.
@@
Materials
Refer to caption Refer to caption
Input
“marble”
“wood”
“clay”
Input
“marble”
“wood”
“clay”
Digital Art
Refer to caption Refer to caption
Input
“3D animation”
“cyberpunk”
“digital painting”
Input
“3D animation”
“cyberpunk”
“digital painting”
Figure 15: Optimization-based editing results for prompts describing different materials and digital art genres. The full prompts are “a marble statue, made from marble, stone”, “a wood sculpture, wooden statue, made from wood”, “clay figure, made from modelling clay”, “a character from a 3D animation movie”, “cyberpunk, futuristic, dystopian”, and “sci-fi, digital painting”.
Famous Artists
Refer to caption Refer to caption
Input
“van Gogh”
“Renoir”
“Vermeer”
Input
“van Gogh”
“Renoir”
“Vermeer”
Art Techniques
Refer to caption Refer to caption
Input
“aquarelle”
“oil paiting”
“pencil sketch”
Input
“aquarelle”
“oil paiting”
“pencil sketch”
Figure 16: Optimization-based editing results for prompts describing artists and art techniques. The full prompts are “art by Vincent van Gogh”, “art by Pierre-Auguste Renoir”, “art by Johannes Vermeer”, “aquarelle painting”, “oil painting”, and “pencil sketch”.

### v2

@@ 5.2 Image-to-image Translation Network Training
Optimization-based editing is not always an optimal choice, e.g. when building interactive editing tools where execution speed matters. Further, optimization can be unstable or perform better or worse, depending on the input images or target prompts. To this end, we explore to what extend LMC-SDS can be used to train a reliable and fast image-to-image translation network. Given a known distribution of images, we train a network to translate images from that distribution to an unseen target distribution, only described by a target caption. Inspired by the authors of DDS [8], we train a cats-to-others network, with “others” being bears, dogs, lions, and squirrels. The network transforms an image of a cat into an image of one of the other species. In DDS, the authors fine-tune a pre-trained diffusion model into a single-step feedforward model using images created by that same diffusion model. In contrast, we train a standard U-Net from scratch and use a dataset of real images, a much more practical and versatile approach. Concretely, we train using cat images from AFHQ [3] and evaluate on 5% held out test images. We use the target prompt “a photo of a [class]” with $\omega=15$ for all methods and the source prompt “a photo of a cat” for DDS. Moreover, we stabilize the training using an L2 loss on the original image and use Cosine scheduling to let the network first focus on reconstructing the original image and then on edits, similarly to the experiment in DDS. The class is passed via a frequency encoded class id concatenated to each pixel. We report numerical results in Tab. 1 and show qualitative examples in Fig. 7. Additional analysis is provided in the Suppl. Mat. SDS performs similarly to previous experiments for this task. While the target class is evident in the resulting images, the results feature strong deficiencies. Again DDS struggles with significant edits and stays close to the source. In contrast, the network trained with our LMC-SDS loss produces realistic and convincing edits. This is also evident from the significantly higher CLIP-R-Precision among all network variants.
@@
Figure 9: Results of DreamFusion using the original SDS loss (right) and our proposed LMC-SDS (left) for $\omega=20$. Our results are much sharper, contain more detail, and feature more realistic colors in all examples.
@@ 5.3 Text-to-3D Asset Generation
In our final experiment, we demonstrate how LMC-SDS can improve text-to-3D models in the style of DreamFusion. In Sec. 3, we discussed how large guidance weights $\omega$ may affect the quality of results. To provide further evidence for this observation, we produce 3D assets with varying $\omega$ using the original DreamFusion formulation and a variant using our LMC-SDS. The results in Fig. 8 also show the trend of high $\omega$ having the tendency to produce oversaturated colors, albeit not as as strongly as in the original paper, possibly due the use of a different image diffusion model. But we also observed an additional issue: the infamous Janus problem, the tendency to repeat detail in each view, becomes more severe with bigger $\omega$, as evident from the eagle example. On the other hand, when using low $\omega$ in the original SDS loss, we observe that DreamFusion struggles to create detailed structure or even structure at all. In contrast, when using LMC-SDS, DreamFusion successfully produces detailed results, even when using up to only one tenth of the originally proposed $\omega$. In Fig. 9 (more in Suppl. Mat.), we show additional side-by-side comparisons of the original SDS-based DreamFusion and our LMC-SDS version. Using our loss, the results are consistently sharper and feature far more detail. Finally in Fig. 10, we show that fixed $\boldsymbol{\epsilon}$ sampling may also be used in 3D asset generation, to obtain diverse results.
@@ 6 Discussion & Conclusion
We proposed a novel loss formulation called LMC-SDS. LMC-SDS allows to effectively utilize text-to-image diffusion models as rich image priors in several optimization and network training tasks. We derived LMC-SDS by analyzing DreamFusion’s SDS loss and by identifying an inherent problem in its formulation. Concretely, we expressed the SDS gradient as the sum of two terms. We identified that the second term $\mathcal{L}_{\text{proj}}$ provides deficient gradients and consequently was weighted down in DreamFusion. We argued that this term, however, should provide gradients towards the learned manifold of natural images associated to the diffusion model. We also provided evidence that the lack thereof makes the optimization unstable, thus negatively impacting the fidelity of results. To this end, we proposed to model the denoising models’ time step dependent frequency bias. This allowed us to factor it out and obtain the cleaner gradient $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{lmc}}$. In extensive experiments, we have demonstrated the effectiveness and versatility of our novel LMC-SDS loss. Furthermore, we discussed how LMC-SDS differs from other SDS-style losses and demonstrated how its performance is superior both qualitatively and quantitatively.
@@
LMC-SDS finds its limits where either the signal from the diffusion model is too weak and ambiguous or when the optimization goes off the natural image manifold by too much. Concretely, LMC-SDS can only utilize whatever signal is coming from the diffusion model. This means, that when the diffusion model does not “understand” the prompt, LMC-SDS (and all other SDS-based approaches) will not be able to provide meaningful gradients. Further, whenever the current optimization state is too corrupted, LMC-SDS will not be able to correct for that and guide the optimization back to the manifold. Also, LMC-SDS sometimes overcompensates the frequency bias. Nevertheless, LMC-SDS is an important first step towards more stable and meaningful usage of the inherent image prior associated to image diffusion models.
We proposed LMC-SDS, a novel SDS-based loss formulation with cleaner gradients towards the image manifold. In the future, we would like to continue to investigate how the manifold corrective can be further improved. We also want to leverage our findings in concrete applications like text-to-3D or in image editing and stylization tasks.
@@
Appendix 0.A Implementation Details
@@
We trained a standard U-Net with 11 layers with two Conv/ReLU/MaxPool blocks per layer, and skip connections from the encoder to the decoder. The first layer contains 32 filters and the number of filters is doubled until 256 in each layer in the encoder and halved in last layers of the decoder. The class id is a float value encoded using 10-dimensional frequency encoding concatenated to each pixel. The final Conv layer with $\tanh$ activation outputs images of $256\times 256$px resolution. We trained the network with a learning rate of $5\times 10^{-5}$ with a decay of $0.9$ every $50$K steps. In the first $20$K steps, we use Cosine scheduling to increase the weight of $\mathcal{L}_{\text{LMC-SDS}}$ from $0.1$ to $1.0$ and decrease the weight of L2 regularization w.r.t. the original image from $1.0$ to $0.1$. We trained for $200$K iterations with batches of $32$ images.
@@
Appendix 0.B Additional Results
Figure 11: Progression of an image optimized using the prompts “photo of a young girl” (top) and “photo of an old woman” (bottom).
@@ Appendix 0.B Additional Results
In Figs. 15, 16, 17 and 18 we present additional editing results. In contrast to the main paper, we show results for each prompt on a number of images. Concretely, we synthesized photographs of people and edited each of them using prompts describing different art styles or materials. Each result respects the corresponding prompt well and yet has an individual touch to it. The image statistics are well preserved and artifacts are rare. The overall quality once more demonstrates the effectiveness of our novel LMC-SDS loss.
@@
Input “marble” “wood” “clay”
Figure 15: Optimization-based editing results for prompts describing different materials. The full prompts are “a marble statue, made from marble, stone”, “a wood sculpture, wooden statue, made from wood”, “clay figure, made from modelling clay”.
Input “3D animation” “cyberpunk” “digital painting”
Figure 16: Optimization-based editing results for prompts describing different digital art genres. The full prompts are “a character from a 3D animation movie”, “cyberpunk, futuristic, dystopian”, and “sci-fi, digital painting”.
Input “van Gogh” “Renoir” “Vermeer”
Figure 17: Optimization-based editing results for prompts describing different artists. The full prompts are “art by Vincent van Gogh”, “art by Pierre-Auguste Renoir”, “art by Johannes Vermeer”.
Input “aquarelle” “oil” “pencil”
Figure 18: Optimization-based editing results for prompts describing different art techniques. The full prompts are “aquarelle painting”, “oil painting”, and “pencil sketch”.